Mr Glocks car broke down on the turnpike.Acme towing charge $30 plus $3 per mile to tow the car.If Mr Glock paid $117 ,how far was the car towed ?
Answer by checkley71(8403)

30+3X=117
3X=117-30
3X=87
X=87/3
X=29 MILES WAS THE DISTANCE HE WAS TOWED.
PROOF
30+3*29=117
30+87=117
117=117